INAF queue observing, 2018 November 06-07

Observer: R. Carini, M. Faccini
Telescope Operator: D. Gonzal-Huerta
Support Astronomer: B. Rothberg
Instrument: LUCI1 (LS mode), LUCI2, LUCI1, MODS1

Summary

Night duration (12\xB0 el twilight) 1:15-12:45 UT = 11h 30m

Observing time: 10h 50min (94%)

Weather time loss: h m (0%)

technical time loss 0h 40m (6%)

Data Summary

At the sunset the sky is clear, wind about 10 m/s, humidity about 60%

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
calibration MODS1 standard mods1 dual grating, flat with slit 5.0"   mods1r.20181106.0068-0070-,mods1b.20181106.0037  
Polletta Telluric HIP106817 luci2 G200 Lores, zjspec   luci2.0011-0014 seeing high
Polletta PLCKG073 luci2 G200 Lores, zjspec 1.23 h luci2.0020-0056 seeing variable, sometimes out of spec
Polletta Telluric HIP106817 luci2 G200 Lores, zjspec   luci2.0060-0063 seeing better
Nardini telluric hip 13917 luci1 G150 Ks, Ks   luci1.0005-0008  
Nardini q2 LBQS 03020019 luci1 G150Ks, Ks 0.72 h luci1.0014-0037 completed
Perna Telluric HIP16111 luci2 G200 LoRes, HKspec, zjspect   luci2.0067-0075  
Perna RCSGA327 luci2 G200LoRes; HKspec, zjspec 0.16 h, 1.12 h luci2.0080-0083, luci2.0084-0111 HKspec completed. seeing variable sometimes out of spec
Prandoni USS1 mods1 red grating, slit 2.4" 0.5 h mods1.0010-0012 seeing variable, sometimes out of spec
standard Feige 34 mods1 red grating   mods1r.0016-0019  
Calibration Prandoni   mods1 flat 2.4"   mods1r.0020-0025  
Calibration Perna   luci2 arc+ flat HKspec,zjspec luci2.0117-0127; luci1.0132-0155  
Calibration Polletta   luci2 arc+flat luci2.0147-0155  
Calibration Severgnini luci2 arc+flat luci2.0161-0165  

Details

Calibraions MODS

flat slit 5.0"
  • blue grating mods1r.20181106.0068-0070-,mods1b.20181106.0037

1:00 AO check

1:32 acq for Polletta Telluric HIP106817
  • luci2.20181107.0008-0010
  • 01:38 sci luci2.0011-0014
  • seeing 2"

01:43 slewing to Polletta PLCKG073
  • allign luci2.0015-0018 luci2.0019
  • seeing 1.0"
  • 2:10 sci luci2.0020-0039
  • 02:19 seeing about 0.7"
  • 02:23 seeing jump to 1.9" and comes back to 1.0"
  • 02:33 seeing variable between 0.9" and 1.2"
  • 02:39 seeing 0.8"
  • 02:48 seeing 1.4"
  • 02:56 seeing about 1.0"
  • 03:03 re send the script to do the second part sci luci2.0040-0056
  • seeing 1.2"
  • 03:17 seeing variable between 1.00" and 1.2"
  • 03:20 seeing 0.9"'1.0"
  • 03:35 seeing jump to 1.6" and comes back to 1.0"-1.2"
  • 03:39 seeing jump to 2.0" on the DIMM (fwhm is 1.3")
  • 03:42 seeing about 1.4" on the DIMM
  • 03:47 the seeing is stable at values more than 1.1" -1.2". we stop the integration

03:49 we take again the telurric since the seeing is very variable HIP106817
  • acq luci2.0057-0059
  • 03:59 sci luci2.0060-0063
  • seeing 1.2"-1.4"

04:03 The seeing is variable around 1.4", we slew to Nisini's program
  • telluric HIP17971
  • 04:05 error MOS

04:24 it will take some times to fix the problem with LUCI2, the AS suggests to slew to LUCI1 or MODS1
  • we decide to use LUCI1, slew to Nardini's program. We wait the AS to start with the telluric.

4:44 sent acq for telluric hip 13917 seeing 1,2" (Nardini's program)

04:49 telluric slewing to hip 13917
  • seeing 1.1"
  • acq luci1.0001-0004
  • 04:54 science luci1.0005-0008

04:59 slew to the target q2
  • acq luci1.0009-0013
  • 05:12 luci1.0014-0037
  • seeing about 1.0"
  • seeing about 1.2"-1.3"

05:29 The support astronomer adivices us that LUCI-2 MOS is recovered * this kind of problem happens when tha mask is changed at low elevation (40 degree (about 1.5 airmass)), the gravity vector affected the robot putting the mask into the cabinet, it couldnt get the angle right. In this case we sould ask to TO to go to zenith first and then we can send the preset for the next target.

06:12 slew to the ZENITH to allow the observation of Perna ob, as suggested by the support astronomer
  • 06:15 send the preset to telluric HIP 16111 acq luci2.0064-000066
  • seeing about 1.0"
  • 06:18 we have to stop to check pointing

06:26 re send preset - seeing about 1.2"

06:36 Sci in filter HK luci2.0067-0070
  • Sci in filter zj luci2.0071-0075

06:48 slewing to Perna science luci2.0080-0083 filter HK

07:39 slewing to Perna science luci2.0084-0107 filter zJ seeing 0.8" *08:09 seeing 1.6" *08:33 re send the script to take some more images luci2.0107-0111

08:55 change instrument for the rest of the night we will use MODS1

09:15 first target with MODS De Carli and send acqMODS Mask22
  • mods1r.0001-0005
  • we have problems with alignment, the masks seems flipped respect the field, the boxes do not correspond to any stars.

10:05 slew to Prandoni program, target USS1
  • problem during the acquisistion, mods can not give the guide coordinate to the telescope.
  • 10:41 fix the problem, in the script there is not the name of the guide star

10:42 send again the acquisition
  • acq mods1.0006-0009
  • sci mods1.0010-0012
  • we start the integration with seeing about 1.0"
  • 11:17 the seeing is increasing, is very variable between 1.2 to 1.8
  • seeing variable between 1.0"-1.3"
  • 11:26 seeing more stble around 1.0''
  • seeing variable between 1.15"and 1.4"
  • seeing 0.8"-1.0"

12:22 slew to standard star, feige34
  • acq mods1r.0013-0014
  • sci =mods1r.0015-0019 ima 0015 no good

Calibrations MODS

Prandoni
  • for the slitless flat, lamps and flat for the standard see Moretti
  • flat with slit 2.4" red grating ==mods1r.0020-0021==controlla se ne ha fatti 1 o 2
  • error :Unknown/Umplemented command STA, we ABort and send again the script.
  • mods1.0022-0025

Calibration Luci

Perna
  • flat with two filters HKspec and zJspec
  • HK_LUCI2.xml
  • flushing luci2.0112-0116
  • for the flats and arc images luci2.0117-0127 HALO2
  • zj_LUCI2.xml flushing luci.0128-0131
  • images luci2.0132-0137 lamp HALO1
  • luci2.0138-0139 lamp NEON
  • luci2.0140-0141 argon

Polletta
  • cal_LUCI2_Fraternal.xml
  • flushing luci2.0142-0146
  • images luci2.0147-0151 lamp HALO2
  • luci2.0152-0153 lamp NEON
  • luci2.0154-0155 lamp ARGON

Severgnini
  • flushing luci2.0156-0160
  • images luci2.0161-0165 lamp HALO1
Topic revision: r9 - 05 Mar 2019, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback